# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v3

@@ Abstract
Advancements in image segmentation play an integral role within the broad scope of Deep Learning-based Computer Vision. Furthermore, their widespread applicability in critical real-world tasks has resulted in challenges related to the reliability of such algorithms. Hence, uncertainty quantification has been extensively studied within this context, enabling the expression of model ignorance (epistemic uncertainty) or data ambiguity (aleatoric uncertainty) to prevent uninformed decision-making. Due to the rapid adoption of Convolutional Neural Network (CNN)-based segmentation models in high-stake applications, a substantial body of research has been published on this very topic, causing its swift expansion into a distinct field. This work provides a comprehensive overview of probabilistic segmentation, by discussing fundamental concepts of uncertainty quantification, governing advancements in the field as well as the application to various tasks. Moreover, literature on both types of uncertainties trace back to four key applications: (1) to quantify statistical inconsistencies in the annotation process due ambiguous images, (2) correlating prediction error with uncertainty, (3) expanding the model hypothesis space for better generalization, and (4) Active Learning. An extensive discussion follows that includes an overview of utilized datasets for each of the applications and evaluation of the available methods. We also highlight challenges related to architectures, uncertainty quantification methods, standardization and benchmarking, and finally end with recommendations for future work such as methods based on single forward passes and models that appropriately leverage volumetric data.
Index Terms:
image segmentation, uncertainty quantification, probabilistic modeling, epistemic, aleatoric
I Introduction
Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest [1]. With the rapid development of Convolution Neural Networks (CNNs), Deep Learning-based image segmentation has seen major advancements and gained significant interest over time [2, 3, 4], obtaining impressive scores with large-scale segmentation datasets [5, 6, 7]. Nevertheless, such methodologies rely on extensive assumptions and relaxations on the Bayesian learning paradigm, omitting crucial information on the uncertainty associated with the model predictions. This omission diminishes the reliability and interpretability of such models. For example, a difficult distinction between classes in real-time automotive scenarios can result in disastrous consequences or uncertain lesion malignancy prediction may significantly impact the decision-making of invasive treatments.
Extensive efforts have been made to align modern neural network optimization with Bayesian Machine Learning [8, 9, 10, 11], such as learning parameter densities, rather than point estimates, to include a notion of epistemic uncertainty. Furthermore, explicitly modeling the output likelihood distribution enables expressing the aleatoric uncertainty. Notably, literature mention that determining the nature of uncertainty is not often straightforward. For example, Hüllermeier et al. [12] stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred”. This sentiment is also shared by Kiureghian and Ditlevsen [13], noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic”. Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic.
Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation and underscore the importance of a rigorous literature overview. Nonetheless, most surveys take the perspective from the medical domain [14, 15], often for specific modalities [16, 17, 18, 19]. There is a notable absence of a comprehensive overview in this field, relating theoretical foundation to the multitude of applications. Furthermore, the abundance of available works can be often overwhelming for both emerging and seasoned researchers. This study seeks to fill this gap in literature and contributes to the subject area by providing a curated overview, where various concepts are discussed through standardized notation and clarified by mutual comparison. After reading this work, readers will be able to discern various forms of uncertainty with their pertinent applications to segmentation tasks. Additionally, they will have developed comprehensive understanding of challenges and unexplored avenues in the field.
This review paper is structured as follows. Past work with significant impact in general image segmentation is presented in Section II, where closely related surveys are also referenced. Then, the theoretical framework and notation that governs the remainder of the paper is introduced in Section III. The role of these concepts in image segmentation will be treated in Sections IV and V, which include all architectures and approaches with significant impact on the field. We then consider the applications that use these uncertainty estimates in Section VI. Finally, this overview will be further discussed together with key challenges and future recommendations in Section VII, which is concluded in Section VIII. Figure 1 illustrates a brief overview of the sections.
@@ I Introduction
II Background
This section highlights a selection of literature crucial to the development of current probabilistic segmentation models. While literature on segmentation itself is vast, this decision has resulted in a limited collection of studies.
As summarized by Minaee et al. [20], semantic segmentation has been performed using methods such as thresholding [21], histogram-based bundling, region-growing [22], k-means clustering [23], watershedding [24], to more advanced algorithms such as active contours [25], graph cuts [26], conditional and Markov random fields [27], and sparsity-based methods [28, 29]. After the application of CNNs [30], the domain of image segmentation underwent rapid developments. Notably, the Fully Convolutional Network (FCN) [3] adapted the AlexNet [31], VGG16 [32] and GoogLeNet [33] architectures to enable end-to-end semantic segmentation. Furthermore, other CNN architectures such as DeepLabv3 [34], and the MobileNetv3 [35] have also been commonly used.
@@ II Background
As the research progressed, increasing success has been observed with encoder-decoder models [36, 4, 37, 2]. Initially developed for the medical applications, Ronneberger et al. [2] introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps. To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures, particularly in the medical domain. In fact, reports of recent research indicate that the relatively simple U-Net and nnU-Net [38] still outperform more contemporary and complex models [39, 40].
Semantic segmentation focuses solely on assigning class labels to each pixel and is particularly suitable for amorphous or uncountable subjects of interest. In contrast, instance segmentation not only detects, but also delineates individual objects within the image. This form of segmentation is more applicable when identifying and outlining countable instances of objects. The third category, panoptic segmentation, combines both class and instance-level classification [20].
III Probabilistic Image Segmentation
This section describes the fundamentals of Bayesian inference and its relationship with conventional models for segmentation. The provided details introduce standard variables and notations as we proceed throughout the manuscript.
III-A Bayesian inference
@@
Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\mathbf{Y},\mathbf{X}}$ that take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image. The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively. Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
@@
As evident, both the variability in the empirical data and the inferred parameters of the model influence the predictive distribution. Hence, uncertainties stemming from the conditional likelihood distribution are classified as either aleatoric, implying from the statistical diversity in the data, or epistemic, which stems from the posterior, i.e. the variance of the model parameters. A straightforward approach to quantify any of these uncertainties is achieved through obtaining the predictive entropy, which can similarly be decomposed as
@@
III-B Conventional segmentation
@@
which simplifies the training procedure by taking a point estimate of the posterior. This approximation improves as the training data increases and the model parameter variances approach zero. As such, MLE does not include any prior knowledge on the structure of the parameter distribution. This is typically achieved through Bayesian Maximum A Posteriori (MAP) estimation with
@@
For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively [41, 42]. Additionally, the output of such models can be interpreted as a probability distribution. For instance, consider spatial image dimensions $D$ and a CNN model $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ with parameters $\bm{\theta}$, such that $\mathbf{a}=f_{\theta}(\mathbf{x})$. Then, the output of such model can be regarded as parameters of a Probability Mass Function (PMF) through
@@
with channel-wise indexing over the denominator, which is commonly known as the SoftMax activation (Figure 2). This approach is probabilistic modeling in the technical sense, although it is not referred to as such in common nomenclature. In fact, the approximated distribution can represent and localize uncertain regions.
However, the implicit pixel-independence assumption
@@
omits information on structural variation in the segmentation masks. At the very core, this limitation (See Figure 3) has driven research on probabilistic segmentation, enabling the sampling of spatially coherent segmentation masks. This challenge can be addressed either from the aleatoric or epistemic perspective. Both approaches have relevant specific use cases, and distinct merits and drawbacks, as will become clear in the upcoming sections.
@@
IV Aleatoric uncertainty
Aleatoric uncertainty quantification reconsiders the non-deterministic relationship between $\mathbf{x}\in\mathcal{X}$ and $\mathbf{y}\in\mathcal{Y}$, which implies that
@@ IV Aleatoric uncertainty
with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$. This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g. not all cats have tails). Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality. The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty.
In most practical scenarios, aleatoric uncertainty quantification methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$. For example, it is possible to model a distribution parameterized by the output of a CNN (Section IV-A). Also, generative models have been used extensively, where the data likelihood is learned through latent variables (Section IV-B). Finally, image augmentation during test-time inference can be also applied to obtain a notion of aleatoric uncertainty (Section IV-C).
IV-A Pixel-level sampling
A valid approach employs direct pixel-level uncertainty in the annotations. In case the independence assumption holds, it is important to ensure proper calibration of the model, which is discussed in Section IV-A1. In another method, the spatial correlation is explicitly modeled across the pixels of the segmentation mask and results in spatially coherent samples, which will be treated in Section IV-A2.
IV-A1 Independence
The normalized confidence values that result from SoftMax activation can only be interpreted as probabilities after proper validation, which is referred to as model calibration. Ideally, the empirical accuracy of a model should be approximately equal to the provided class confidence $c_{k}$ for class $k$, i.e. $P(Y=k\,|\,c_{k})=c_{k}$. Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over-confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4). Different methods can be used to measure calibration, but often introduce their own biases. A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy (acc) and confidence (conf) bins as
with $n_{b}$ the number of samples in bin $b$ and $N$ being the total sample size across all bins. The ECE is prone to skew representations if some bins are significantly more populated with samples due to over-/under-confidence. Furthermore, the Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.

### v4

@@ Abstract
Advances in architectural design, data availability, and compute have driven remarkable progress in semantic segmentation. Yet, these models often rely on relaxed Bayesian assumptions, omitting critical uncertainty information needed for robust decision-making. The resulting reliance on point estimates has fueled interest in probabilistic segmentation, but the literature remains fragmented, marked by inconsistent terminology, evaluation practices, and theoretical ambiguities. In response, this review consolidates and contextualizes foundational concepts in uncertainty modeling, including the non-trivial task of distinguishing between epistemic and aleatoric uncertainty and examining their roles across four key downstream segmentation tasks, highlighting Active Learning as particularly promising. By unifying theory, terminology, and applications, we provide a coherent foundation for researchers and identify critical challenges, such as strong assumptions in spatial aggregation, lack of standardized benchmarks, and pitfalls in current uncertainty quantification methods. We identify trends such as the adoption of contemporary generative models, driven by advances in the broader field of generative modeling, with segmentation-specific innovation primarily in the conditioning mechanisms. Moreover, we observe growing interest in distribution- and sampling-free approaches to uncertainty estimation. We further propose directions for advancing uncertainty-aware segmentation in deep learning, including pragmatic strategies for disentangling different sources of uncertainty, novel uncertainty modeling approaches and improved Transformer-based backbones. In this way, we aim to support the development of more reliable, efficient, and interpretable segmentation models that effectively incorporate uncertainty into real-world applications.
1 Introduction
Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest (Szeliski, 2010). The advent of convolutional neural networks (CNNs) has led to major breakthroughs in this domain, with deep learning-based methods achieving state-of-the-art performance on large-scale datasets (Ronneberger et al., 2015; Shelhamer et al., 2014; Badrinarayanan et al., 2015), obtaining impressive scores with large-scale segmentation datasets (Lin et al., 2014; Cordts et al., 2016b; Richter et al., 2016). However, these models typically rely on strong assumptions and significant relaxations of the Bayesian learning paradigm, neglecting the uncertainty associated with their predictions. This lack of uncertainty modeling reduces both the reliability and interpretability of the predictions. In high-stakes applications, such as autonomous driving or medical diagnosis, this can have severe consequences. For instance, misclassifying adjacent objects in autonomous driving or overlooking uncertainty in lesion classification can both lead to critical decision-making errors.
Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation, especially as interpretability and reliability have become central in data-driven applications. Extensive efforts have been made to align neural network optimization with Bayesian machine learning (Blundell et al., 2015; Guo et al., 2017; Kingma & Welling, 2013; Kendall & Gal, 2017), such as learning parameter distributions rather than point estimates to capture epistemic uncertainty. Additionally, explicitly modeling the likelihood of outputs enables estimation of aleatoric uncertainty. However, in practice, these two forms of uncertainty are often entangled or even conflicting (Mucsányi et al., 2024; de Jong et al., 2024; Valdenegro-Toro & Mori, 2022; Hüllermeier & Waegeman, 2021), making the distinction difficult to define and apply. This conceptual ambiguity presents significant barriers to understanding and implementation (Kahl et al., 2024; Der Kiureghian & Ditlevsen, 2009; Kirchhof et al., 2025). At the same time, uncertainty is often treated as an auxiliary tool to enhance downstream performance, rather than as a primary modeling objective. As a result, many works lack rigorous theoretical grounding, focusing instead on empirical gains without clear definitions or principled formulations of uncertainty. Furthermore, many approaches are tailored to specific datasets or modalities, limiting generalizability. Moreover, theoretical contributions often originate from adjacent fields like Bayesian deep learning or information theory, without being properly contextualized in segmentation. This disconnect has resulted in a fragmented body of work, with inconsistent terminology, evaluation metrics, and assumptions. The resulting abundance of literature can be overwhelming, even for experienced researchers.
This paper aims to explore and clarify these concepts in the context of segmentation architectures and their impact on downstream tasks. Given these challenges and the renewed interest in uncertainty (Papamarkou et al., 2024; Kahl et al., 2024; Kirchhof et al., 2025), we identify a notable absence of a comprehensive, theory-informed overview in the field. Due to its application-driven nature, most existing surveys adopt a medical perspective (Jungo & Reyes, 2019; Kwon et al., 2020), often focused on specific modalities (McCrindle et al., 2021; Jungo et al., 2020; Ng et al., 2022; Roshanzamir et al., 2023). However, the connection between theoretical foundations and their diverse applications remains underexplored. Additionally, the study by Kahl et al. (2024) introduces a valuable framework for benchmarking uncertainty disentanglement in semantic segmentation, but focuses primarily on empirical evaluation within specific tasks. In contrast, our work provides a comprehensive overview that spans the full range of established uncertainty quantification methods, unifies and contextualizes the underlying theory, and standardizes terminology and notation. This broader perspective helps clarify conceptual foundations and highlights cross-cutting insights across segmentation architectures and application domains. By the end of this paper, readers will have a clear understanding of the various forms of uncertainty, their relevance to segmentation tasks, and a comprehensive grasp of the key challenges and open research directions in the field.
@@ 1 Introduction
This review paper is structured as follows. Past work with significant impact in general image segmentation is presented in Section 2. Then, the theoretical framework and notation that govern the remainder of the paper is introduced in Section 3. The theory enables us to classify the approaches into two distinct but often related methods. That is, aleatoric and epistemic uncertainty modeling methods. The role of these concepts in image segmentation will be treated in Sections 4 and 5, respectively, which include all architectures and approaches with significant impact on the field. We then correlate these modeling methods to specific applications in Section 6. Here, we also observe for which domains uncertainty quantification has been exploited. Then, this overview will be discussed further in Section 7. This section is particularly valuable for researchers seeking to identify the most suitable model for their specific segmentation task. To encourage further research in this field, we provide future recommendations by highlighting key challenges, gaps and pitfalls in literature, concluding our review in Section 8. Figure 1 illustrates a brief overview of the sections.
2 Background
A common approach in image understanding involves labeling pixels according to semantic categories, which is the core of semantic segmentation. This technique is particularly well-suited for amorphous or uncountable subjects. In contrast, instance segmentation not only assigns class labels, but also distinguishes and delineates individual object occurrences. This makes it more appropriate for scenarios involving countable entities. A third variant, panoptic segmentation, unifies both semantic and instance-level labeling, offering a comprehensive view of scene composition. As summarized by Minaee et al. (2020), semantic segmentation has been performed using methods such as thresholding (Otsu, 1979), histogram-based bundling, region-growing (Dhanachandra et al., 2015), k-means clustering (Nock & Nielsen, 2004), watershedding (Najman & Schmitt, 1994), to more advanced algorithms such as active contours (Kass et al., 2004), graph cuts (Boykov et al., 2001), conditional and Markov random fields (Plath et al., 2009), and sparsity-based methods (Starck et al., 2005; Minaee & Wang, 2017). While the literature on segmentation is vast and rapidly evolving, a selection of backbone architectures has been particularly influential in shaping current probabilistic segmentation models. The focus here is not on exhaustiveness, but on those models most relevant to the development of uncertainty-aware approaches.
@@ 2 Background
In particular, following the successful application of CNNs (LeCun et al., 1998), image segmentation experienced rapid progress driven by increasingly powerful and specialized deep architectures. Notably, the Fully Convolutional Network (FCN) (Shelhamer et al., 2014) adapted the AlexNet (Krizhevsky et al., 2012), VGG16 (Simonyan & Zisserman, 2014) and GoogLeNet (Szegedy et al., 2014) architectures to enable end-to-end semantic segmentation. Furthermore, other CNN architectures such as DeepLabv3 (Chen et al., 2017), and the MobileNetv3 (Howard et al., 2019) have also been commonly used. As the research progressed, increasing success has been observed with encoder-decoder models (Noh et al., 2015; Badrinarayanan et al., 2015; Yuan et al., 2019; Ronneberger et al., 2015). Initially developed for the medical applications, Ronneberger Ronneberger et al. (2015) introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps. To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures, particularly in the medical domain. In fact, reports of recent research indicate that the relatively simple U-Net and nnU-Net (Isensee et al., 2020) still outperform more contemporary and complex models (Eisenmann et al., 2023; Isensee et al., 2024).
3 Probabilistic Image Segmentation
@@
Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\*Y,\*X}$ that take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image. The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively. Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
@@
As evident, both the variability in the empirical data and the inferred parameters of the model influence the predictive distribution. Hence, uncertainties stemming from the conditional likelihood distribution are classified as either aleatoric, implying from the statistical diversity in the data, or epistemic, which stems from the posterior, i.e. the variance of the model parameters. Encapsulating a particular uncertainty can therefore be achieved by incorporating parametrized stochasticity at either feature- or parameter-level. A straightforward approach to quantify any of these uncertainties is achieved by obtaining the predictive entropy, $H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]$, i.e. the entropy of the predictive distribution $p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})$. Furthermore, disentangling these uncertainties can be achieved through its decomposition
@@
In this way, it is theoretically possible to identify whether high-entropy predictions are due to model ignorance or statistical ambiguity inherent in the data generating process. Nonetheless, determining the nature of uncertainty is not often straightforward. For example, Hüllermeier & Waegeman (2021) stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred". This sentiment is also shared by Der Kiureghian & Ditlevsen (2009), noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic". Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic as we treat the realm of quantifying spatially correlated uncertainty. Especially with increasingly complex methodologies, treating the uncertainties as separate concepts is mostly theoretical (often even more philosophical) and highly non-trivial in practice.
3.1 Conventional segmentation
@@
which simplifies the training procedure by taking a point estimate of the posterior. This approximation improves as the training data increases and the model parameter variances approach zero. As such, MLE does not include any prior knowledge on the structure of the parameter distribution but can be achieved through Bayesian Maximum A Posteriori (MAP) estimation with
@@
For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively (Figueiredo, 2001; Kaban, 2007). Additionally, the output of such models can be interpreted as a probability distribution. For instance, consider spatial image dimensions $D$ and a CNN model $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ with parameters $\bm{\theta}$, such that $\mathbf{a}=f_{\theta}(\mathbf{x})$. Then, the output of such model can be regarded as parameters of a Probability Mass Function (PMF) through
@@
with channel-wise indexing over the denominator, which is commonly known as the SoftMax activation (Figure 2). This approach is probabilistic modeling in the technical sense, although it is not referred to as such in common nomenclature. In fact, the approximated distribution can represent and localize uncertain regions. However, the implicit pixel-independence assumption
@@
omits information on structural variation in the segmentation masks. In conventional classification, factorizing the categorical distribution is typically regarded as a logical simplification. In probabilistic segmentation however, this assumption has caused the emergence of a distinct research direction (See Figure 3). The challenge of sampling of spatially coherent segmentation masks can be addressed either from the aleatoric or epistemic perspective, usually indicating reducability of the assumed uncertainty source. Both approaches have specific use cases, and each modeling choice comes with distinct advantages and limitations, which will be discussed in the following sections.
@@
4 Aleatoric uncertainty
[…]
@@ 4 Aleatoric uncertainty
with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$. This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g. not all cats have tails). Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality. The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty. In most practical scenarios, aleatoric uncertainty quantification methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$. For example, it is possible to model a distribution parameterized by the output of a CNN (Section 4.1). Also, conditional generative models are used to learn the data distribution through latent variables. For the purpose of taxonomy, we categorize models based on the location of these latent (unobserved) variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional latents embedded deeper within the architecture are covered in Section 4.2. While their theoretical formulations are largely similar, their practical implications can differ significantly.
4.1 Pixel-level sampling
Uncertainty in segmentation masks can be modeled directly at the pixel level. These approaches can be further categorized into those that assume independence between pixels (Section 4.1.1), and those that explicitly model spatial correlations (Section 4.1.2). In the former case, accurate uncertainty estimates rely on well-calibrated models, which is an assumption that often fails in practice without additional tuning on a separate calibration set. In the latter case, a stochastic variable is introduced to capture dependencies between neighboring pixels.
4.1.1 Independence